OpenStudy (anonymous):

8 sqrt 7s)/(sqrt 28s^2) Sorry doing this from my phone is a little hard to type.

OpenStudy (anonymous):

Simplify

OpenStudy (anonymous):

I got 4s

OpenStudy (harsha19111999):

Can you tell by giving spaces? I couldn't understand

OpenStudy (anonymous):

\[8\sqrt{7s}/\sqrt{28s ^{2}}\]\[(8\sqrt{7s}*\sqrt{28s ^{2}})/28s ^{2}\]\[(8\sqrt{7s}*\sqrt{s ^{2}}*\sqrt{28})/28s ^{2}\]\[(8s \sqrt{7}*\sqrt{s}*\sqrt{4}*\sqrt{7})/28s ^{2}\]\[(16s \sqrt{s}\sqrt{7*7})/28s ^{2}\]\[(112s \sqrt{s})/28s ^{2}\]\[(4s \sqrt{s})/s ^{2}\]\[(4\sqrt{s})/s\]